OpenStudy (anonymous):

Please do help me! I'm very confused!! Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

@ParthKohli @campbell_st @thomaster @beccaboo333 @wio @eliassaab @timo86m @wolfe8 @hba @shubhamsrg @Xmoses1 @tester97

OpenStudy (anonymous):

From the point (2, -2), 2 = rcos(θ) and -2 = rsin(θ). Dividing the two equations, \[\frac{2}{-2}=\frac{rcos(θ)}{rsin(θ)}\]which gives you\[-1=tan(θ)\]Can you solve this equation to get the value of θ?

OpenStudy (anonymous):

@navk 1/4 is what I'm getting...

OpenStudy (anonymous):

The equation above \[-1=tan(\theta)\]is supposed to give you \[\theta=-pi/4\]

OpenStudy (anonymous):

Oh-- Well, I got 1/4 from wolframalpha..

OpenStudy (anonymous):

So, it would be 45 and 225 degrees, right?

OpenStudy (anonymous):

it is -45 degrees or 225 degrees, because tan(theta) is negative

OpenStudy (anonymous):

(2 square root of 2, 225°), (-2 square root of 2, 45°) (2 square root of 2, 45°), (-2 square root of 2, 225°) Which one out of these two then?

OpenStudy (anonymous):

Plugging the value of theta in the equation we get \[2=rcos(-pi/4)\]which gives you\[2=r*1/sqrt(2)\]that is\[r=2sqrt(2)\]hence the coordinates are (2sqrt(2), -45 degrees) or (2sqrt(2), 225 degrees)

OpenStudy (anonymous):

thank you so much, have a nice day!! c: